Find the measure of an angle such that the difference between its supplement and 2 times its complement is 33 degrees
Answer by mananth(16949)

let angle be x
supplement = 180-x
complement =90-x
(180-x)-2(90-x)=33
180-x -180+2x=33
x=33
the angle = 33